OpenStudy (eric_d):

the expression ax^3-8x^2+bx+6 is divisible by x^2-2x-3.Find the values of a and b.

OpenStudy (eric_d):

@ShailKumar

OpenStudy (anonymous):

Can you try to divide ?

OpenStudy (eric_d):

hw.. there are unknowns

ganeshie8 (ganeshie8):

if it is divisible by x^2-2x-3, then it will be divisible by its individual factors : x-3 and x+1 also since x^2-2x-3 = (x-3)(x+1)

ganeshie8 (ganeshie8):

by remainder/factor theorem, x-3 is a factor => f(3) = 0 x+1 is a factor => f(-1) = 0 use them to solve a, b

OpenStudy (eric_d):

I need to form two equations?

OpenStudy (mathstudent55):

\(ax^3-8x^2+bx+6\) \( = (mx + n)(x^2-2x-3)\) \(= mx^3 - 2 mx^2 - 3mx + nx^2 - 2nx - 3n \) \(= mx^3 + (-2m + n)x^2 + (-3m - 2n)x - 3n \) -3n = 6 n = -2 -3m - 2n = b -3m - 2(-2) = b -3m + 4 = b Eq. 1 -2m + n = - 8 -2m + (-2) = -8 -2m = -6 m = 3 Eq. 1: -3m + 4 = b -3(3) + 4 = b -9 + 4 = b b = -5 a = m a = 3

OpenStudy (mathstudent55):

A colorful explanation of the lines above. Line 1: \(\color{red}{a}x^3+ \color{blue}{-8}x^2+\color{green}{b}x+\color{brown}{6}\) Line 4: \(= \color{red}{m}x^3 + \color{blue}{(-2m + n)}x^2 + \color{green}{(-3m - 2n)}x + \color{brown}{- 3n}\)

OpenStudy (eric_d):

Hw you form the second line.. (mx+n)(x^2-2x-3) @mathstudent55

OpenStudy (mathstudent55):

Since you are told a third degree polynomial is divisible by a second degree polynomial, then the quotient must be a first degree polynomial, which we know is a binomial of the form mx + n.

OpenStudy (eric_d):

Ok..

OpenStudy (eric_d):

Thanks @mathstudent55 @ganeshie8 @ShailKumar

OpenStudy (mathstudent55):

You're welcome.